Establishing the North East Greenland Ice Stream (NEGIS) camp – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

NEEM > Field diaries > Field season 2012 > Establishing the North...

13 June 2012

Establishing the North East Greenland Ice Stream (NEGIS) camp

Today’s mud and ice core from below the ice sheet.

A final Polar 6 shuttle brought four scientists to the remote NEGIS site some 400 km SE of NEEM where a camp is now established. The scientists will stay at NEGIS for four weeks doing seismic work and radar surveying until they are picked up by Polar 6 in July. During this period, NEEM camp will be in contact with NEGIS on a daily basis.

The shallow ice core drilling at NEGIS was terminated at 66.1 m depth and the ice core and the shallow drill are now back at NEEM.

What we have done today:
1. Drilling sediment cores with rock drill in the deep hole
2. Polar 6 flight to NEGIS bringing crew in and drill out
3. Established camp at NEGIS
4. Shallow drilling at NEGIS: 61 m – 66.1 m depth
5. DEPping NEGIS core. Laki found at 44.3 m depth
6. Setting up equipment in the water vapor sampling tent
7. Shooting film all over camp for Euronews
8. Drilling sample holes in skiway and in balloon cave roof
9. Preparing supporting beams in garage for lifting operation
10.Marking new foam boxes using numbers 5994-5999

Ad. 1: Today penetration went slowly. First run resulted in 13 cm core (see picture) while a second longer run was left at the bottom.

Ad. 3: The 4 person NEGIS camp crew was deployed and a camp was established. A first contact was made on Iridium (phone # in NEEM field log book page 1). Everyone is in good shape. NEGIS camp will call NEEM every morning at 8 am. If connection fails, they continue calling every half hour.

Ad. 4: The drilling at NEGIS continued and was terminated at 66.1 m depth.
Due to technical issues with the drill it was not possible to proceed and all equipment was brought back to NEEM.

Ad. 6: Motohiro is setting up equipment in the water vapor sampling tent.
A small AWS is installed and a proton transfer reaction mass spectrometer for DMS detection was unpacked. Unfortunately, a vital stainless steel tube had broken during transport and the instrument is unlikely to function until a spare part has been brought to camp.

Ad. 8: The skiway was probed by the hand auger for 3D tomography microstructure analysis at AWI.

Weather: Another beautiful day with clear blue sky and fog banks in the morning. At night a cloud cover rolled in. Temperatures  –9°C to – 20°C, wind 5-12 knots from SW slowly picking up.

FL, Anders Svensson


← Previous entry                                                                                 Next entry →